# Condensing and Extracting Against Online Adversaries

### v2

@@ 3.1 Basic Probability Notions
Lemma 3.2 (Min-entropy chain rule, [maurer1997privacy]).
@@ 3.2 Condensers and Extractors
Seeded extractor is the special case of seeded condenser where $k_{out}=m$. We here record the full definition for completeness sake:
@@
Theorem 3.5 (Theorem 1.5 in [guruswami_unbalanced_2009]).
For all constant $10<\alpha<$ 1, there exists a constant $C$ such that for all $n,k,\varepsilon$, there exists an explicit $(k,\varepsilon)$-seeded extractor $\mathsf{sExt}:\{0,1\}^{n}\times\{0,1\}^{d}\rightarrow\{0,1\}^{m}$ with $d=C\log(n/\varepsilon)$ and $m\geq(1-\alpha)k$.
@@
3.3 Leader Election, Collective Coin Flipping and Sampling Protocols
We formalize the definition of protocols in the full information model. Collective coin flipping protocols, leader election protocols, and collective sampling protocols are special cases of such protocols where the output domain is $[\ell]$ and $\{0,1\}$ and $\{0,1\}^{m}$ for some $m$ respectively.
@@ 3.3 Leader Election, Collective Coin Flipping and Sampling Protocols
Here is how the protocol operates in the presence of a set $B\subset[\ell]$ of bad players: In round $i$, each of the players from $[\ell]\setminus B$ independently output a uniformly random element from $\{0,1\}^{n}$. Let their collective outputs be $\alpha_{i}\in\left(\{0,1\}^{n}\right)^{[\ell]\setminus B}$. Then, depending on $\alpha_{1},\dots,\alpha_{i}$, the players in $B$ together output an element of ($\left(\{0,1\}^{n}\right)^{B}$. Hence, we model the strategy of the bad players as a sequence of functions $\sigma=(\sigma_{1},\dots,\sigma_{k})$, where

### v3

@@ 4.1 Basic Probability Notions
We also state the useful folklore result of the data processing inequality.
Fact 4.2.
For any two distributions $\mathbf{X},\mathbf{Y}$ over $\Omega$ and function $f:\Omega\to\mathbb{R}$,
$\left\lvert\mathbf{X}-\mathbf{Y}\right\rvert\geq\left\lvert f(\mathbf{X})-f(\mathbf{Y})\right\rvert$.
@@
Lemma 4.3 (Min-entropy chain rule, [MW97]).
@@ 4.2 Condensers and Extractors
A seeded extractor is the special case of seeded condenser where $k_{out}=m$. We record the full definition for completeness sake:
@@
Theorem 4.6 (Theorem 1.5 in [GUV09]).
For all constant $0<\alpha<1$, there exists a constant $C$ such that for all $n,k,\varepsilon$, there exists an explicit $(k,\varepsilon)$-seeded extractor $\mathsf{sExt}:\{0,1\}^{n}\times\{0,1\}^{d}\rightarrow\{0,1\}^{m}$ with $d=C\log(n/\varepsilon)$ and $m\geq(1-\alpha)k$.
@@
4.3 Averaging Samplers
Recall the definition of an averaging sampler
Definition 4.8.
A $(k,\delta,\varepsilon)$-averaging sampler is a function $\mathsf{Samp}:\{0,1\}^{n}\to(\{0,1\}^{m})^{D}$ such that for any function $f:\{0,1\}^{m}\to[0,1]$ and any $(n,k)$-source $\mathbf{X}$, we have that
$\Pr_{(x_{1},\dots,x_{D})\sim\mathsf{Samp}(\mathbf{X})}\left[\left\lvert\frac{1}{D}\sum_{i=1}^{D}f(x_{i})-\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{m}}[f(x)]\right\rvert\geq\varepsilon\right]\leq\delta.$
It was shown in [Zuc97] that strong extractors and averaging samplers are equivalent. We reproduce the proof here for completeness.
Lemma 4.9.
Let $\mathsf{Ext}:\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m}$ be a $(k,\varepsilon)$-extractor. Define $\mathsf{Samp}:\{0,1\}^{n}\to(\{0,1\}^{m})^{D}$, where $D=2^{d}$ as $\mathsf{Samp}(x)=(\mathsf{Ext}(x,1),\dots,\mathsf{Ext}(x,D))$ where we identify $[D]$ with $\{0,1\}^{d}$. Then $\mathsf{Samp}$ is a $(k+\log(1/\delta),\delta,\varepsilon)$-averaging sampler.
Proof.
Let $\mathbf{X}$ be an $(n,k+\log(1/\delta))$-source. For $x\in\operatorname{Supp}(\mathbf{X})$, call $x$ bad if $\left\lvert\mathsf{Ext}(x,\mathbf{U}_{d})-\mathbf{U}_{m}\right\rvert>\varepsilon$ and let $B\subseteq\{0,1\}^{n}$ be the subset of bad $x$’s. Now, suppose for the sake of contradiction that $\Pr[\mathbf{X}\in B]>\delta$. Then, letting $\mathbf{X}_{B}=\mathbf{X}\mid(\mathbf{X}\in B)$, we can compute the min-entropy of $\mathbf{X}_{B}$ as
$H_{\infty}(\mathbf{X}_{B})$ $\geq H_{\infty}(\mathbf{X})-\log\left(\frac{1}{\Pr[\mathbf{X}\in B]}\right)$
$\geq k+\log(1/\delta)-\log(1/\delta)$
$=k.$
Therefore, we can apply $\mathsf{Ext}$ to $\mathbf{X}_{B}$ and obtain that $\left\lvert\mathsf{Ext}(\mathbf{X}_{B},\mathbf{U}_{d})-\mathbf{U}_{n}\right\rvert\leq\varepsilon$. However, by assumption we know that for all $x\in B$, $\left\lvert\mathsf{Ext}(x,\mathbf{U}_{d})-\mathbf{U}_{m}\right\rvert>\varepsilon$, meaning that $\left\lvert\mathsf{Ext}(\mathbf{X}_{b},\mathbf{U}_{d})-\mathbf{U}_{m}\right\rvert>\varepsilon$, giving us a contradiction. Thus, we have that $\Pr[\mathbf{X}\in B]\leq\delta$.
Now, we turn our attention to the good $x\notin B$. Using 4.2, we know that for any $x\notin B$,
[…]
$\leq\varepsilon.$
This is equivalent to saying that, for all good $x\in\{0,1\}^{n}$, $\left\lvert\frac{1}{D}\sum_{s\in\{0,1\}^{d}}f(\mathsf{Ext}(x,s))-f(\mathbf{U}_{m})\right\rvert\leq\varepsilon$. This is exactly the requirement of our sampler, and we have shown that this happens with probability $\Pr[\mathbf{X}\notin B]\geq 1-\delta$, as required. ∎
In particular, we will use the following strong extractor from [Zuc07] to instantiate an averaging sampler.
Theorem 4.10 ([Zuc07]).
For all constant $\alpha,\delta,\varepsilon>0$, there is an efficient family of strong $(k=\delta n,\varepsilon)$-extractors $\mathsf{Ext}:\{0,1\}^{n}\times\{0,1\}^{d}\to\{0,1\}^{m}$ with $m\leq(1-\alpha)\delta n$ and $D=2^{d}=O(n)$.
Using Lemma 4.9, we get the following averaging sampler that we shall use later.
Lemma 4.11.
For all constant $\alpha,\delta,\varepsilon>0$, we can construct an explicit sampler $\mathsf{Samp}:\{0,1\}^{t}\to(\{0,1\}^{m})^{D}$ with $m\leq(1-\alpha)\delta t$ and $D=O(t)$ such that for all sets $S\subseteq[M]$, where $M=2^{m}$, and all $(t,k)$-sources $X$, we have that
$\Pr_{x\sim X}\left[\left\lvert\frac{\left\lvert\mathsf{Samp}(x)\cap S\right\rvert}{D}-\frac{\left\lvert S\right\rvert}{M}\right\rvert\geq\varepsilon\right]\leq 2^{\delta t-k}.$
Proof.
Simply apply Lemma 4.9 to Theorem 4.10 and consider the indicator function $f(x)=1_{x\in S}$ of $S$ for the resulting sampler. ∎
4.4 Leader Election, Collective Coin Flipping, and Sampling Protocols
We formalize the definition of protocols in the full information model. Collective coin flipping protocols, leader election protocols, and collective sampling protocols are special cases of such protocols where the output domain is $\ell$] and $\{0,1\}$ and $\{0,1\}^{m}$ for some $m$ respectively.
@@ 4.4 Leader Election, Collective Coin Flipping, and Sampling Protocols
Here is how the protocol operates in the presence of a set $B\subset[\ell]$ of bad players: In round $i$, each of the players from $\ell]\setminus B$ independently output a uniformly random element from $\{0,1\}^{n}$. Let their collective outputs be $\alpha_{i}\in\left(\{0,1\}^{n}\right)^{[\ell]\setminus B}$. Then, depending on $\alpha_{1},\dots,\alpha_{i}$, the players in $B$ together output an element of $\left(\{0,1\}^{n}\right)^{B}$. Hence, we model the strategy of the bad players as a sequence of functions $\sigma=(\sigma_{1},\dots,\sigma_{k})$, where